1120ml of ozonised oxygen at s.t.p weigh 1.76g.calculate the volume of oxygen in the ozonised oxygen.

Grade:12th pass

4 Answers

AKSHAT TECHNICALS
6 years ago
note that, molecular mass of oxygen = 32 molecular mass of ozone = 48 calculations:- let vol of O2 = x ml hence vol of O3 = 1120 - x ml at STP, wt of x ml O2 = 32*x/22400 gm wt of 1120 - x ml O3 = 48*(1120 - x)/22400 gm hence, 32*x/22400 + 48*(1120 - x)/22400 = 1.76 or, 32x + 53760 - 48x = 39424 => 16x = 53760-39424=14336=> x = 14336/16=896ml
AKSHAT TECHNICALS
6 years ago
note that, 
molecular mass of oxygen = 32 
molecular mass of ozone = 48 
 
calculations:- 
 
let vol of O2 = x ml 
hence vol of O3 = 1120 - x ml 
at STP, 
wt of x ml O2 = 32*x/22400 gm 
wt of 1120 - x ml O3 = 48*(1120 - x)/22400 gm 
 
hence, 
32*x/22400 + 48*(1120 - x)/22400 = 1.76 
or, 32x + 53760 - 48x = 39424 
=> 16x = 53760-39424=14336
  • => x = 14336/16=896ml
k sitaram Rao
one year ago
A sample of ozonised oxygen contains 90% by volume of oxygen  what will be volume of resulting gases and hen 120 ml of such sample 
k sitaram Rao
one year ago
A sample of ozonised oxygen contains 90% by
volume of oxygen. What will be the volume of
the resulting gases when 120ml of such a sample
is heated and then cooled to the original conditions
of temperature and pressure ?
1) 126 ml 2) 132 ml 3) 100 ml
4) 50 ml